# Hybrid Memetic Search for Electric Vehicle Routing with Time Windows, Simultaneous Pickup-Delivery, and Partial Recharges

### v1

@@ V-B Results on Small-scale and Medium-scale Instances
HMA w/o BCD HMA
Instance m best avg±std time m best avg±std time p-value
jd200_1 44 70740.90 71335.98±403.87 3678.82 44 70440.57 71409.68±512.27 3510.36 0.6501
jd200_2 39 59389.70 60004.86±419.68 3661.48 38 58530.22 59823.10±570.84 3612.10 0.3643
jd200_3 42 65201.82 65844.88±445.08 3544.28 43 65812.71 66206.45±284.52 3534.71 0.0588
jd200_4 39 60747.28 60983.75±250.76 3571.75 39 60815.01 61008.17±184.95 3362.58 0.4497
jd400_1 78 114537.34 116518.79±1422.92 10492.89 78 114489.28 115579.17±715.05 10474.83 0.1306
jd400_2 76 115303.21 116803.18±612.69 10409.36 77 114668.85 115658.43±606.08 10815.96 0.0015
jd400_3 76 106951.59 107842.05±588.35 10547.36 76 107600.70 107927.90±351.95 10063.53 0.8206
jd400_4 81 126613.16 127573.34±914.21 10687.95 82 126927.71 127814.74±492.32 10362.35 0.2265
jd600_1 112 166524.71 168729.52±1110.96 14626.36 112 166113.91 166952.86±681.25 18016.97 0.0019
jd600_2 113 171996.76 173256.93±1387.16 14450.07 113 170362.24 171446.42±618.81 17416.13 0.0007
jd600_3 110 151879.40 153745.61±1118.40 16444.28 109 151185.33 152554.96±942.57 17806.93 0.0494
jd600_4 103 149893.97 151118.39±658.62 13495.48 103 148729.43 149447.53±563.44 17432.51 0.0004
jd800_1 147 216194.15 217442.46±805.04 20030.45 146 213065.47 214228.94±974.71 25217.34 0.0002
jd800_2 148 226218.76 227630.94±1710.11 19306.58 147 222333.51 223685.24±1338.04 25214.00 0.0007
jd800_3 147 205429.64 207285.21±1202.21 23965.29 146 203751.04 205483.36±1355.49 24410.31 0.0082
jd800_4 135 194738.56 196802.01±952.89 23059.46 134 192848.13 194144.22±987.70 24949.64 0.0005
jd1000_1 178 260502.17 262307.88±1199.71 29046.07 178 256626.16 257932.03±1114.15 32159.10 0.0002
jd1000_2 182 275972.37 278824.43±1564.13 28589.01 181 270788.78 273195.12±1537.57 32421.83 0.0002
jd1000_3 184 256229.79 257553.76±1262.58 25397.05 181 250802.03 253040.02±1411.33 31452.40 0.0002
jd1000_4 169 242662.53 244305.00±1250.99 27539.25 167 238562.78 239966.65±1244.47 32276.48 0.0002
@@ V-C Results on Large-scale Instances
To demonstrate the effectiveness of HMA for solving large-scale instances in the $jd$ set, we limited both short and long maximum computation time and compared HMA with HMA without the BCD decomposition (denoted HMA w/o BCD). Here, we omit the existing algorithms [11, 17] for solving EVRP-TW-SPD because they are not open-sourced. Specifically, instances with 200, 400, 600, 800, and 1000 customers were decomposed into 2, 4, 6, 8, and 10 subproblems in the MS of HMA (see Algorithm 3). The short time limits were set to 1800, 5400, 9000, 12600, and 16200 seconds (see Table IV), while the long time limits were 3600, 10800, 18000, 25200, and 32400 seconds (see Table V). Note that BCD requires the Cartesian coordinate of node $n_{i}$, while the $jd$ set contains longitude and latitude. Since the area covered is not large, we can use Mercator Projection to convert $(lng_{i},lat_{i})$ into $(x_{i},y_{i})$, specifically, $x_{i}=R\cdot\frac{\pi\cdot lng_{i}}{180}$, $y_{i}=R\cdot\ln\left(\tan\left(\frac{\pi}{4}+\frac{\pi\cdot lat_{i}}{360}\right)\right)$, where $R$ is the semi-major axis of the WGS84 ellipsoid model. In Tables IV and V, the “avg±std” columns represent the average and standard deviation of $TC$ for the 10 best solutions found in 10 independent runs. We used the Wilcoxon rank-sum test with significance level $p=0.05$ to compare the performance of HMA w/o BCD and HMA, with p-values shown in the last column. The best performance w.r.t. “best” is highlighted in gray, and w.r.t. “avg±std” is highlighted in bold if it is significantly better than the other algorithm, respectively.
@@
From Table IV, it can be observed that under a short time limit, HMA consistently found better solutions on all instances of the $jd$ set compared to HMA w/o BCD. The average performance of HMA is significantly better as the problem scale increases. In Table V, under a long time limit, HMA w/o BCD performed reversely better on some instances with 200 and 400 customers. This is due to the challenge of designing an ideal decomposition, which would partition all customers belonging to the same route in the optimal solution into the same subproblem. However, other decompositions, such as BCD, solve subproblems efficiently but may result in suboptimal solutions. Nonetheless, as the problem scale grows, the decomposition strategy still shows an obvious advantage on instances of larger scale with 600, 800, and 1000 customers since HMA obtained significantly higher-quality solutions. The performance gap between HMA w/o BCD and HMA across 10 independent runs, measured as $[TC_{1}-TC_{2}]/{TC}_{2}$, where $TC_{1}$ and $TC_{2}$ are the value of best-found solutions from HMA w/o BCD and HMA, is 0.61%, 1.26%, and 1.83% averaged on instances with 600, 800, and 1000 customers, respectively. As the problem scale grows, the performance gap between HMA w/o BCD and HMA also becomes larger.
Figure 2: Comparative effects of PSI and SSI on each subset. The outer ring represents the exclusive contribution ratio (ECR) of SSI and PSI, with the remainder being shared contributions; The inner ring shows the time allocation ratio (TAR) of SSI and PSI, with the remainder representing others’ TAR.
V-D Effectiveness of PSSI, CDNS, and Hybrid Framework
@@
First, we separately analyzed the comparative effects of PSI and SSI for the following reasons: PSSI selects the best result from PSI and SSI. Therefore, we are more concerned about which strategy, PSI or SSI, achieves better station insertion on a route. Note that SSI produces only a single route, compared to PSI with a population. We need to validate whether PSI can really find other higher-quality station insertion configurations with acceptable additional computational cost. To this end, we ran HMA once on all instances, keeping track of the performance of PSI and SSI in terms of station insertion on the same route. Figure 2 presents the performance of PSI and SSI: the exclusive contribution ratio (ECR) is computed as $n_{1}/n_{2}$ for each instance, where $n_{1}$ is the number of insertion results obtained exclusively from SSI or PSI and $n_{2}$ is the total number of insertion results; the time allocation ratio (TAR) is computed as $t_{1}/t_{2}$ for each instance, where $t_{1}$ is the cumulative time consumed in SSI or PSI components, $t_{2}$ is the total time consumed in HMA. Note that the outer ring “ECR” and inner ring “TAR” in Figure 2 are the average values on each subset.
@@
From the results, it can be clearly seen that PSI can really find other higher-quality station insertions, as ECR $>0$ indicates that it found some strictly better insertions than SSI. In the $jd$ set, the ECR of PSI even achieved 46.90%, accounting for a large proportion of contributions to PSSI. Notably, although PSI with a genetic algorithm took approximately $2\sim 8$ times runtime than SSI, this additional runtime is worthwhile and acceptable, as SSI cannot further improve its insertion results. Thus, combining the results of both PSI and SSI can better avoid local optima and construct higher-quality routes with acceptable additional computational cost.
@@
Second, we conducted an ablation study to evaluate the effectiveness of other components in HMA. We tested the following four variants on the $akb$ and $jd$ sets, each of which is different from HMA in one component:
@@
    w/o CLS: lines 2-2 of Algorithm 2 are removed;
@@
    w/o ALS: lines 2-2 of Algorithm 2 are removed;
@@
    w/o MS: lines 3-3 of Algorithm 3 are removed;
@@
    w/o LNS: lines 3-3, 3-3 of Algorithm 3 are removed.
The first two variants were to validate the effectiveness of CDNS, while the latter two were to validate the effectiveness of the hybrid framework. For each variant, the performance degradation ratio (PDR) is computed as $[TC_{1}-TC_{2}]/{TC}_{2}$ for each instance, where $TC_{1}$ and $TC_{2}$ are the average $TC$ obtained by the variant and HMA across 10 independent runs, respectively. PDR $>0$ indicates the component is useful for this specific instance. Larger values indicate a greater impact. Table VI presents the average PDR on each subset.

### v2

@@ V-B Results on Small-scale and Medium-scale Instances
Adapt-CMSA-STD Adapt-CMSA-SETCOV HMA w/o BCD HMA
Instance name best avg±$\pm$±std time $\Delta\%$ best avg±$\pm$±std time $\Delta\%$ best avg±$\pm$±std time $\Delta\%$ best avg±$\pm$±std time p-value
jd200_1 73644.20 74971.07$\pm$892.62 3601.47 76125.24 78512.42$\pm$1482.09 3609.07 70740.90 71335.98$\pm$403.87 3678.82 70440.57 71409.68$\pm$512.27 3510.36 0.6501
jd200_2 65007.91 66759.93$\pm$1100.01 3577.74 71075.99 71940.92$\pm$728.62 3605.35 59389.70 60004.86$\pm$419.68 3661.48 58530.22 59823.10$\pm$570.84 3612.10 0.3643
jd200_3 67644.15 69089.36$\pm$946.59 3196.51 71600.11 72482.85$\pm$646.16 3604.60 65201.82 65844.88$\pm$445.08 3544.28 65812.71 66206.45$\pm$284.52 3534.71 0.0588
jd200_4 71929.75 74205.32$\pm$1681.07 3601.88 9.17% 81696.93 84631.25$\pm$2017.07 3609.60 18.16% 60747.28 60983.75$\pm$250.76 3571.75 0.21% 60815.01 61008.17$\pm$184.95 3362.58 0.4497
jd400_1 134302.39 137821.08$\pm$1868.47 10804.12 143045.76 146142.97$\pm$1767.66 10822.74 114537.34 116518.79$\pm$1422.92 10492.89 114489.28 115579.17$\pm$715.05 10474.83 0.1306
jd400_2 134084.89 135903.71$\pm$1220.10 11032.40 138520.94 141625.25$\pm$1845.82 10813.94 115303.21 116803.18$\pm$612.69 10409.36 114668.85 115658.43$\pm$606.08 10815.96 0.0015
jd400_3 115586.81 116862.96$\pm$798.97 10434.63 120065.47 123138.26$\pm$1663.56 10734.00 106951.59 107842.05$\pm$588.35 10547.36 107600.70 107927.90$\pm$351.95 10063.53 0.8206
jd400_4 135410.22 137420.59$\pm$1633.07 10744.00 12.09% 137863.84 141825.04$\pm$1917.25 10813.89 16.49% 126613.16 127573.34$\pm$914.21 10687.95 -0.06% 126927.71 127814.74$\pm$492.32 10362.35 0.2265
jd600_1 202030.04 207678.74$\pm$4499.67 18498.28 210658.62 215812.21$\pm$2820.95 17241.05 166524.71 168729.52$\pm$1110.96 14626.36 166113.91 166952.86$\pm$681.25 18016.97 0.0019
jd600_2 201625.22 205362.93$\pm$2455.99 18544.71 205943.56 211244.27$\pm$3406.55 16875.98 171996.76 173256.93$\pm$1387.16 14450.07 170362.24 171446.42$\pm$618.81 17416.13 0.0007
jd600_3 169213.04 171977.52$\pm$2543.89 18403.92 173106.97 176575.91$\pm$2006.10 15701.68 151879.40 153745.61$\pm$1118.40 16444.28 151185.33 152554.96$\pm$942.57 17806.93 0.0494
jd600_4 188018.90 191041.29$\pm$2450.90 18016.40 19.58% 197929.76 201306.90$\pm$2408.88 18013.41 23.82% 149893.97 151118.39$\pm$658.62 13495.48 0.61% 148729.43 149447.53$\pm$563.44 17432.51 0.0004
jd800_1 264947.80 271723.12$\pm$8551.65 25368.63 270951.52 280903.46$\pm$9941.59 23268.07 216194.15 217442.46$\pm$805.04 20030.45 213065.47 214228.94$\pm$974.71 25217.34 0.0002
jd800_2 263423.50 269372.58$\pm$3767.75 25209.97 268047.17 278252.06$\pm$12695.67 25214.92 226218.76 227630.94$\pm$1710.11 19306.58 222333.51 223685.24$\pm$1338.04 25214.00 0.0007
jd800_3 226143.04 232084.76$\pm$2892.31 26087.54 232052.94 244010.43$\pm$15964.86 21723.70 205429.64 207285.21$\pm$1202.21 23965.29 203751.04 205483.36$\pm$1355.49 24410.31 0.0082
jd800_4 248967.19 260163.60$\pm$9966.13 25228.11 20.73% 260642.67 267303.51$\pm$7378.45 25225.15 24.19% 194738.56 196802.01$\pm$952.89 23059.46 1.26% 192848.13 194144.22$\pm$987.70 24949.64 0.0005
jd1000_1 323977.19 337199.82$\pm$10774.46 32435.92 325780.40 329586.90$\pm$2529.39 32469.22 260502.17 262307.88$\pm$1199.71 29046.07 256626.16 257932.03$\pm$1114.15 32159.10 0.0002
jd1000_2 332121.12 341578.59$\pm$8647.74 32909.71 332438.38 335741.59$\pm$2112.72 32477.35 275972.37 278824.43$\pm$1564.13 28589.01 270788.78 273195.12$\pm$1537.57 32421.83 0.0002
jd1000_3 283186.67 288387.04$\pm$2503.00 33144.56 286423.55 290715.99$\pm$2296.68 32463.47 256229.79 257553.76$\pm$1262.58 25397.05 250802.03 253040.02$\pm$1411.33 31452.40 0.0002
jd1000_4 316816.36 323327.80$\pm$5526.63 32451.29 23.65% 320716.02 324125.48$\pm$3133.73 32407.54 24.59% 242662.53 244305.00$\pm$1250.99 27539.25 1.83% 238562.78 239966.65$\pm$1244.47 32276.48 0.0002
@@ V-C Results on Large-scale Instances
To demonstrate the effectiveness of HMA for solving large-scale instances in the $jd$ set, we compared HMA with three algorithms: Adapt-CMSA-STD [13], Adapt-CMSA-SETCOV [19], and HMA without the BCD decomposition (denoted HMA w/o BCD). Since the original implementations of Adapt-CMSA-STD and Adapt-CMSA-SETCOV are not made open-source, we implemented them based on the descriptions provided in the literature, ensuring high fidelity to the reported algorithms. The BCD decomposition strategy integrated into HMA is tailored for large-scale problem instances as follows. Specifically, instances with 200, 400, 600, 800, and 1000 customers were decomposed into 2, 4, 6, 8, and 10 subproblems, respectively, during the MS of HMA (see Algorithm 3). To evaluate the impact of runtime on algorithm performance for large-scale instances, both short and long time limits were imposed, i.e., the short time limits were set to 1800, 5400, 9000, 12600, and 16200 seconds (see Table IV), while the long time limits were 3600, 10800, 18000, 25200, and 32400 seconds (see Table V). Note that the BCD requires the Cartesian coordinates of node $n_{i}$, whereas the $jd$ set contains the longitude and latitude values. Since the covered area is relatively small, we can use the Mercator projection to convert the geographic coordinates $(lng_{i},lat_{i})$ into Cartesian coordinates $(x_{i},y_{i})$. Specifically, the conversion is given by $x_{i}=R\cdot\frac{\pi\cdot lng_{i}}{180}$, $y_{i}=R\cdot\ln\left(\tan\left(\frac{\pi}{4}+\frac{\pi\cdot lat_{i}}{360}\right)\right)$, where $R$ is the semi-major axis of the WGS84 ellipsoid model. Tables IV and V present the testing results of the four algorithms on large-scale instances in the $jd$ set. Unlike the $akb$ set, for the $jd$ set there is no priority for minimizing either the number of used EVs or $TD$, thus the “m” values are not reported, but the “best” values are still included. The “avg±$\pm$±std” columns represent the average and standard deviation of $TC$ for the 10 best solutions obtained from 10 independent runs. The performance gap “$\Delta\%$” is measured as $[TC_{1}-TC_{2}]/{TC}_{2}$, where $TC_{1}$ and $TC_{2}$ represent the best-found solutions from the other three competitors and HMA over 10 independent runs. Here we show the average “$\Delta\%$” values for each problem size. The best performance with respect to “best” is highlighted in gray. The following analysis focuses on two comparisons: i) the two HMA variants (i.e., HMA and HMA w/o BCD) against the two Adapt-CMSA variants (i.e., Adapt-CMSA-STD and Adapt-CMSA-SETCOV), and ii) HMA against HMA w/o BCD.
V-C1 HMA Variants versus Adapt-CMSA Variants
The results in Tables IV and V clearly show that the two HMA variants consistently achieved superior performance over the two Adapt-CMSA variants in terms of both the best solution found and the average solution quality across 10 independent runs. A closer examination of the standard deviations for the two Adapt-CMSA variants reveals their unstable performance on large-scale instances, indicating they may struggle to escape poor-quality local optima. In contrast, the two HMA variants exhibit smaller standard deviations across all instances, suggesting a more reliable search process and stability. The performance gap of the two Adapt-CMSA variants can be observed from the “$\Delta\%$” columns. As the problem size increases, this “$\Delta\%$” value also increases under both short and long time limits, highlighting HMA’s remarkable advantage in larger-scale instances. This trend implies that, although Adapt-CMSA-STD and Adapt-CMSA-SETCOV may be adequate for smaller problems, they do not scale as effectively as HMA when confronted with larger instances.
The scalability of HMA can be attributed to the following components. First, when tackling large-scale instances, the large neighborhood search of the HMA framework plays a crucial role, as it enables efficient exploration of the solution space. Second, the aggressive local search is indeed effective. From the analysis of solutions obtained by HMA, we observe that the utilization rate of charging stations is relatively low, and each route typically serves a small number of customers. This characteristic aligns well with practical scenarios. It ensures that ALS can improve solutions in the VRP-TW-SPD solution space without being prematurely terminated due to the early occurrence of PSSI’s failure attempt (see Algorithm 2, line 2, the second termination condition). Moreover, this characteristic guarantees that the binary sequences initialized in PSI do not result in many infeasible solutions (see the ablation study on the $jd$ set in Section V-E).
V-C2 HMA versus HMA w/o BCD
@@
We further evaluated the impact of BCD by comparing HMA w/o BCD and HMA using the Wilcoxon rank-sum test at a 0.05 significance level (see the last column in Tables IV and V). This comparison was necessary because the performance difference between the two HMA variants was not obvious (both average “$\Delta\%$” values on all instances are less than 1.00% under different time limits). The best performance with respect to “avg±$\pm$±std” is highlighted in bold, provided it is statistically significantly better than the other algorithm. From Table IV, it can be observed that under a short time limit, HMA consistently found better solutions on all instances of the $jd$ set compared to HMA w/o BCD. The average performance of HMA is significantly better as the problem scale increases. In Table V, under a long time limit, HMA w/o BCD performed reversely better on some instances with 200 and 400 customers. This is due to the challenge of designing an ideal decomposition, which would partition all customers belonging to the same route in the optimal solution into the same subproblem. However, other decompositions, such as BCD, solve subproblems efficiently but may result in suboptimal solutions. Nonetheless, as the problem scale grows, the decomposition strategy still shows an obvious advantage on instances of larger scale with 600, 800, and 1000 customers since HMA obtained significantly higher-quality solutions. The performance gap “$\Delta\%$” between HMA w/o BCD and HMA across 10 independent runs is 0.61%, 1.26%, and 1.83% averaged on instances with 600, 800, and 1000 customers, respectively. As the problem scale grows, this performance gap becomes even larger, confirming the effectiveness of the BCD decomposition in helping HMA solve very large EVRP-TW-SPD instances.
Overall, these experiments confirm that i) the proposed HMA framework provides the best and robust performance across all large-scale instances in the $jd$ set, and ii) the BCD decomposition is particularly beneficial when tackling instances with 600 or more customers. Even with extended runtimes, HMA consistently outperforms the two Adapt-CMSA variants and yields significantly better solutions than its non-BCD counterpart on the instances with 400 or more customers (see Section V-D for the statistical test). Consequently, the scalability of HMA makes it well-suited for real-world EVRP-TW-SPD applications in which problem instances can easily grow to several hundred or even a thousand customers.
(a) all instances
(b) small-scale instances
(c) medium-scale instances
(d) large-scale instances, a short time limit
(e) large-scale instances, a long time limit
(f) clustered instances
(g) random instances
(h) random-clustered instances
(i) type-1 instances
(j) type-2 instances
(k) 200-customer instances
(l) 400-customer instances
(m) 600-customer instances
(n) 800-customer instances
(o) 1000-customer instances
Figure 2: Critical difference (CD) plots across various problem instances using the Nemenyi post hoc test.
V-D Statistical Test of Algorithm Performance and Robustness
To support our claim that HMA has a significant performance advantage over existing methods across a wide range of problem instances, we also present critical difference (CD) plots. These plots serve as a statistical tool to help evaluate the obtained results in Tables II, III, IV and V, and to measure the algorithm’s robustness when comparing multiple algorithms over various problem instances [19].
The procedure of the statistical test is described as follows. We first apply the Friedman test to simultaneously compare all algorithms and determine if significant differences exist in their rankings. If the null hypothesis, which assumes equal performance among the algorithms, is rejected, we proceed with pairwise comparisons using the Nemenyi post hoc test. The results are then graphically displayed in CD plots (see Figure 2), where each algorithm is positioned along the horizontal axis according to its average rank over the specific subset of problem instances. In total, we consider fifteen subsets. Sub-figures 2(a)2(e) illustrate various problem scales, Sub-figures 2(f)2(j) show different customer distributions and constraint characteristics (see Section V-A1), and Sub-figures 2(k)2(o) provide a more in-depth analysis of large-scale scenarios. Algorithms that exhibit performance differences within the critical difference threshold, computed at a significance level of 0.05, are considered statistically equivalent, as indicated by horizontal lines connecting their markers.
As shown in Figure 2, HMA is statistically significantly better than its competitors in almost all subsets, except for the small-scale instances. In this subset, Adapt-CMSA-STD, Adapt-CMSA-SETCOV, and HMA show comparable performance, with no significant differences. In Sub-figures 2(a)2(e), Adapt-CMSA-SETCOV outperforms Adapt-CMSA-STD on medium-scale instances; however, in large-scale instances with a long time limit, the results are reversed. HMA performs better than HMA w/o BCD in large-scale instances under different time limits. In Sub-figures 2(f)2(j), HMA demonstrates greater strength on random, random-clustered, and type-1 instances. Adapt-CMSA-SETCOV is the runner-up on $akb$ set; In Sub-figures 2(k)2(o), HMA and HMA w/o BCD perform similarly in instances with 200 customers, but a significant difference emerges as the scale increases from 400 to 1000 customers. Adapt-CMSA-STD performs better than Adapt-CMSA-SETCOV with 200 and 400 customers; however, in instances with 1000 customers, the results are reversed. Overall, these CD plots demonstrate that HMA exhibits the best and the most robust performance across different instance scales, customer distributions, constraint characteristics, and time limits.
V-E Effectiveness of PSSI, CDNS, and Hybrid Framework
@@
First, we separately analyzed the comparative effects of PSI and SSI for the following reasons: PSSI selects the best result from PSI and SSI. Therefore, we are more concerned about which strategy, PSI or SSI, achieves better station insertion on a route. Note that unlike PSI, which maintains a population of routes, SSI produces only a single route. We need to validate whether PSI can indeed find other higher-quality station insertion configurations with acceptable additional computational cost. To this end, we ran HMA once on all instances, tracking the performance of PSI and SSI in terms of station insertion on the same route. Figure 3 illustrates this performance. The exclusive contribution ratio (ECR) is calculated as $n_{1}/n_{2}$ for each instance, where $n_{1}$ is the number of insertion results obtained exclusively from SSI or PSI and $n_{2}$ is the total number of insertion results. The time allocation ratio (TAR) is computed as $t_{1}/t_{2}$ for each instance, where $t_{1}$ is the cumulative time consumed in SSI or PSI components, and $t_{2}$ is the total time consumed in HMA. Note that the left bar labeled “ECR” and the right bar labeled “TAR” in Figure 3 represent the average values for each subset.
@@
The results clearly indicate that PSI can indeed find other higher-quality station insertions, as an ECR greater than 0 shows that it found some insertions that are strictly better than those produced by SSI. In the $jd$ set, PSI achieved an ECR of 46.90%, accounting for a substantial portion of PSSI’s overall performance. Notably, although PSI with a genetic algorithm required approximately 22228888 times the runtime of SSI, this additional time is worthwhile and acceptable, since SSI cannot further improve its insertion results. Thus, combining the results of both PSI and SSI can better avoid local optima and construct higher-quality routes with acceptable additional computational cost.
Figure 3: Comparative effects of PSI and SSI on each subset. The left stacked bar illustrates the exclusive contribution ratio (ECR) of SSI and PSI, with the remaining portion corresponding to shared contributions. The right stacked bar depicts the time allocation ratio (TAR) of SSI and PSI, with the remainder representing the TAR of other algorithm components.
@@
Second, we conducted an ablation study to evaluate the effectiveness of other components in HMA. We tested the following four HMA variants on the $akb$ and $jd$ sets, each of which is different from HMA in one component:
@@
    w/o CLS: lines 22 of Algorithm 2 are removed;
@@
    w/o ALS: lines 22 of Algorithm 2 are removed;
@@
    w/o MS: lines 33 of Algorithm 3 are removed;
@@
    w/o LNS: lines 33, 33 of Algorithm 3 are removed.
The first two variants were to validate the effectiveness of CDNS, while the latter two were to validate the effectiveness of the hybrid framework. For each variant, the performance degradation ratio (PDR) is computed as $[TC_{1}-TC_{2}]/{TC}_{2}$ for each instance, where $TC_{1}$ and $TC_{2}$ are the average $TC$ obtained by the variant and HMA across 10 independent runs, respectively. A PDR greater than 0 indicates the component is useful for this specific instance. Larger PDR values indicate a greater impact. Table VI presents the average PDR on each subset.